# Optical-vortex-pulse induced nonequilibrium spin textures in spin-orbit coupled electrons

### v1

@@ III.3 ${\rm SU}(2)$ symmetric SOI: $\beta_{\rm D}=\alpha_{\rm R}$
Let us now examine the spin response in the case of $\alpha_{\rm R}=\beta_{\rm D}$, which significantly differs from that in the case of purely Rashba-type or Dresselhaus-type SOIs. Figure 4 shows the linear response of the local spin to vortex beams with $(m,\lambda)=(1,1)$ and $(2,1)$. Similarly to the case of purely Rashba-type SOIs, the vortex beam causes spin polarization in the region where the intensity of the vortex beams peaks. However, as one sees from Fig. 4, the photo-induced spin moment predominantly points to one direction, i.e., $[1\bar{1}0]$ direction, reflecting the symmetry at $\alpha_{\rm R}=\beta_{\rm D}$. In Fig. 5, we also plot the snapshots of the spin response to the pulsed vortex beam. At all times, the responded spins align to the $[1\bar{1}0]$ direction, and their intensity reflects a vortex shape influenced by the pulsed light. The movies demonstrating the real-time spin dynamics of 2DEGs with $\alpha_{\rm R}k_{\rm F}=\beta_{\rm D}k_{\rm F}=0.1~{}{\rm meV}$ are available in Supplemental Materilal [82].
@@
[…]
where $E_{+}\equiv(E_{x}+E_{y})/\sqrt{2}$. The numerical results shown in Figs. 4 and 5 align with Eq. (22), where all the spins are oriented to the $[1\bar{1}0]$ direction at all times.
@@ III.4 Orbital angular momentum dependence of spin response
Figure 6: Maximal spin polarization, $\max|S({\bm{x}},t)/S_{0}|$, as a function of the OAM of light, $m\in[-6,6]$. The red (blue) circles represent the results for $\beta_{\rm D}/\alpha_{\rm R}=0$ ($\beta_{\rm D}/\alpha_{\rm R}=1$). Here we set $\alpha_{\rm R}k_{\rm F}=0.1~{}{\rm meV}$ and $T/\varepsilon_{\rm F}=0.1$.
@@
Lastly, we examine the OAM dependencies of the optical spin response in spin-orbit coupled electron gases. In Fig. 6, we plot the maximal spin polarization, $\max|{\bm{S}}({\bm{x}},t)|/S_{0}$, as a function of the OAM of light, $m\in[-6,6]$. Here, $\max|{\bm{S}}({\bm{x}},t)|$ means the maximum value of $|{\bm{S}}({\bm{x}},t)|$ over the entire range of $({\bm{x}},t)$. In the case of the Rashba-type SOI ($\beta_{\rm D}/\alpha_{\rm R}=0$), the conventional Gaussiam beam, corresponding to optical vortices with $m=0$, causes nonvanishing spin polarization and the maximal spin polarization gradually increases with $|m|$. In contrast, the spins do not respond to the Gaussian beams when $\alpha_{\rm R}=\beta_{\rm D}$. This is consistent with the previous work [83], where the spins of the 2DEG at the ${\rm SU}(2)$ symmetric point do not respond to the applied DC field, resulting in zero spin polarization.
@@
It is also worth mentioning that for Gaussian beams, the maximal spin polarization at $\alpha_{\rm R}=\beta_{\rm D}$ is several orders of magnitude smaller compared to that in the case of $\beta_{\rm D}/\alpha_{\rm R}=0$. This is attributed to the ${\rm SU}(2)$ symmetry emerging at $\alpha_{\rm R}=\beta_{\rm D}$, where the effective magnetic field ${\bm{b}}_{\rm eff}({\bm{k}})=\alpha_{\rm R}(k_{x}+k_{y})(1,-1,0)$, defined in Eq. (3), causes unidirectional spin alignment. The intensity profile of conventional Gaussian beams peaks around $q=0$ in the momentum space, implying that the spin response is governed by the spatially uniform component of the spin-current response function, i.e., $S_{\mu}({\bm{x}},\omega)\approx\Upsilon_{\mu\nu}({\bm{q}}={\bm{0}},\omega)E_{\nu}({\bm{q}},\omega)$. On the other hand, an optical vortex with nonzero $m$ is a structured light, and its intensity profile is a doughnut-like shape in the transverse plane. The spin response of 2DEGs peaks in the doughnut-shaped region, and the radius of the doughnut increases with $|m|$. For the vortex beam with a larger $|m|$, the contributions of the larger $|{\bm{q}}|$ to the spin-current correlation functions become significant, leading to the strong OAM dependence of the spin polarization. As shown in Fig. 6, the maximal spin polarization at $\alpha_{\rm R}=\beta_{\rm D}$ increases with increasing $|m|$ and reaches $0.002~{}\mu{\rm m}^{-2}$ for $E_{0}=0.1~{}{\rm kV}/{\rm cm}$. These results demonstrate that vortex beams enable ultrafast optical manipulation of spins even at the ${\rm SU}(2)$ symmetric point $\alpha_{\rm R}=\beta_{\rm D}$ for the realistic values of the AC electric field, $E_{0}$.
@@ IV Conclusion
Finally, we comment on a possible pump-probe experimental method of detecting the photo-induced spin textures. The in-plane spin textures can be detected using the longitudinal or transverse magneto-optical Kerr effect [84, 85]. A linearly polarized probe beam is incident at an angle to the normal direction of the 2DEGs, and the reflected light captures information about the in-plane magnetization through the longitudinal or transverse Kerr rotation.
@@ Acknowledgements.
We thank Ken Morita for useful comments on the detection of in-plane spin textures. This work was supported by JST CREST (Grant No. JPMJCR24R5), JSPS KAKENHI (Grants No. JP23K20828, No. JP23K22492, No. JP25H00599, No. JP25H00609, No. JP25K07227, No. JP25K07198), a Grant-in-Aid for Transformative Research Areas (A) “Correlation Design Science” (Grant No. JP25H01250 and No. JP25H01251), “Evolution of Chiral Materials Science using Helical Light Fields” (Grant No. JP22H05131, JP23H04576 and JP25H01609), and “Chimera Quasiparticles for Novel Condensed-Matter Science” (Grant No. JP25H02112) from JSPS of Japan.
@@ Appendix A Impurity vertex correction
Figure 7: Diagrammatic representation of ladder-type impurity scattering for vertex corrections ($\Lambda_{\mu}$) in spin-current response functions. The thick curves denote the Green’s function $G(k)$, and the dashed lines connected to the symbol “$\times$” represent impurity scattering.
@@
In the main text, however, we consider the limit of weak impurity concentration, $\omega\tau\sim\Omega\tau\gg 1$. In this regime, the function, $\Pi_{00}$ can be evaluated by using $G^{\rm R}_{s}({\bm{k}},\varepsilon)=P\frac{1}{\varepsilon-E_{s}}-i\pi\delta(\varepsilon-E_{s})$ for $\Gamma\rightarrow 0$, where $P$ denotes the Cauchy principal part. Substituting this into Eq. (36), one finds ${\rm Re}\Pi_{00}=0$ and $-{\rm Im}\Pi_{00}\propto\frac{1}{\omega}$. This indicates that in the limit of $\omega\tau\gg 1$, the spin vertex function reduces to the bare spin operator, $\hat{\Lambda}_{\mu}\approx\hat{S}_{\mu}$. While we have considered Rashba-type SOIs, the same conclusion is valid for ${\rm SU}(2)$ symmetric SOIs as well.

### v2

@@ III.3 ${\rm SU}(2)$ symmetric SOI: $\beta_{\rm D}=\alpha_{\rm R}$
Let us now examine the spin response in the case of $\alpha_{\rm R}=\beta_{\rm D}$, which significantly differs from that in the case of purely Rashba-type or Dresselhaus-type SOIs. Figure 4 shows the linear response of the local spin to vortex beams with $(m,\lambda)=(1,1)$ and $(2,1)$. Similarly to the case of purely Rashba-type SOIs, the vortex beam causes spin polarization in the region where the intensity of the vortex beams peaks. However, as one sees from Fig. 4, the photo-induced spin moment predominantly points to one direction, i.e., $1\bar{1}0]$ direction, reflecting the symmetry at $\alpha_{\rm R}=\beta_{\rm D}$. In Fig. 5, we also plot the snapshots of the spin response to the pulsed vortex beam. At all times, the responded spins align to the $1\bar{1}0]$ direction, and their intensity reflects a vortex shape influenced by the pulsed light. The movies demonstrating the real-time spin dynamics of 2DEGs with $\alpha_{\rm R}k_{\rm F}=\beta_{\rm D}k_{\rm F}=0.1~{\rm meV}$ are available in Supplemental Materilal [82].
@@
$S({\bm{x}},t)=\int dt^{\prime}\int d{\bm{x}}^{\prime}\tilde{\Upsilon}_{z+}({\bm{x}}-{\bm{x}}^{\prime},t-t^{\prime})E_{+}({\bm{x}}^{\prime},t^{\prime}),$ (25)
where $E_{+}\equiv(E_{x}+E_{y})/\sqrt{2}$. The numerical results shown in Figs. 4 and 5 align with Eq. (24), where all the spins are oriented to the $1\bar{1}0]$ direction at all times.
@@ III.4 Orbital angular momentum dependence of spin response
Figure 6: Maximal value of local spin polarization, $\max|S({\bm{x}},t)/S_{0}|$, as a function of the OAM of light, $m\in[-6,6]$. The red (blue) circles represent the results for $\beta_{\rm D}/\alpha_{\rm R}=0$ ($\beta_{\rm D}/\alpha_{\rm R}=1$). Here we set $\alpha_{\rm R}k_{\rm F}=0.1~{\rm meV}$ and $T/\varepsilon_{\rm F}=0.1$. We note that the total spin is zero, $\int{\bm{S}}({\bm{x}},t)d{\bm{x}}={\bm{0}}$, for $\alpha_{\rm R}=\beta_{\rm D}$.
@@
Lastly, we examine the OAM dependencies of the optical spin response in spin-orbit coupled electron gases. In Fig. 6, we plot the numerically computed maximal spin polarization, $\max|{\bm{S}}({\bm{x}},t)|/S_{0}$, as a function of the OAM of light, $m\in[-6,6]$. Here, $\max|{\bm{S}}({\bm{x}},t)|$ means the maximum value of the laser-driven local spin magnitude $|{\bm{S}}({\bm{x}},t)|$ over the entire range of $({\bm{x}},t)$. In the case of the Rashba-type SOI ($\beta_{\rm D}/\alpha_{\rm R}=0$), the conventional Gaussiam beam, corresponding to optical vortices with $m=0$, causes nonvanishing spin polarization and the maximal spin polarization gradually increases with $|m|$. In contrast, the spins do not respond to the Gaussian beams when $\alpha_{\rm R}=\beta_{\rm D}$. This is consistent with the previous work [83], where the spins of the 2DEG at the ${\rm SU}(2)$ symmetric point do not respond to the applied DC field, resulting in zero spin polarization.
@@
It is also worth mentioning that for Gaussian beams, the maximal spin polarization at $\alpha_{\rm R}=\beta_{\rm D}$ is several orders of magnitude smaller compared to that in the case of $\beta_{\rm D}/\alpha_{\rm R}=0$. This is attributed to the ${\rm SU}(2)$ symmetry emerging at $\alpha_{\rm R}=\beta_{\rm D}$, where the effective magnetic field ${\bm{b}}_{\rm eff}({\bm{k}})=\alpha_{\rm R}(k_{x}+k_{y})(1,-1,0)$, defined in Eq. (3), causes unidirectional spin alignment. The intensity profile of conventional Gaussian beams peaks around $q=0$ in the momentum space, implying that the spin response is governed by the spatially uniform component of the spin-current response function, i.e., $S_{\mu}({\bm{x}},\omega)\approx\Upsilon_{\mu\nu}({\bm{q}}={\bm{0}},\omega)E_{\nu}({\bm{q}},\omega)$. On the other hand, an optical vortex with nonzero $m$ is a structured light, and its intensity profile is a doughnut-like shape in the transverse plane. The spin response of 2DEGs peaks in the doughnut-shaped region, and the radius of the doughnut increases with $|m|$. For the vortex beam with a larger $|m|$, the contributions of the larger $|{\bm{q}}|$ to the spin-current correlation functions become significant. As shown in Figs. 4 and 5, vortex beams with nonzero $m$ can induce a locally nontrivial spin texture while keeping the total spin at zero, i.e.,
$\int{\bm{S}}({\bm{x}},t)d{\bm{x}}=0,$ (26)
under spin-helix conditions, $\alpha_{\rm R}=\beta_{\rm D}$. This results in the strong OAM dependence of the local spin polarization. As shown in Fig. 6, the maximal spin polarization at $\alpha_{\rm R}=\beta_{\rm D}$ increases with increasing $|m|$ and reaches $0.002~\mu{\rm m}^{-2}$ for $E_{0}=0.1~{\rm kV}/{\rm cm}$. These results demonstrate that vortex beams enable ultrafast optical manipulation of spins even at the ${\rm SU}(2)$ symmetric point $\alpha_{\rm R}=\beta_{\rm D}$ for the realistic values of the AC electric field, $E_{0}$.
@@ IV Conclusion
In this paper, the numerical calculations rely on linear response theory to aim at capturing the central aspect of spin manipulation using vortex beams. However, the retardation effect caused by the dipolar coupling of spins and a more precise treatment of impurity effects–beyond the weak impurity approximation–may play an important role when discussing spin dynamics in a realistic situation. These effects fall outside the scope of our current theory and remain as potential future work.
Finally, we comment on a possible pump-probe experimental method of detecting the photo-induced spin textures. The in-plane spin textures can be detected using the longitudinal or transverse magneto-optical Kerr effect [84, 85]. A linearly polarized probe beam is incident at an angle to the normal direction of the 2DEGs, and the reflected light captures information about the in-plane magnetization through the longitudinal Kerr rotation. Using the magneto-optical Kerr microscopy, the magnetization and magnetic domains in a material with in-plane anisotropy has been be obtained [86]. Another way is to induce the precession mode of in-plane electron spins by applying a static magnetic field. For example, it has recently been demonstrated that the higher-order spatial structure of photon polarization can be coherently transferred to an electron system in GaAs/AlGaAs quantum wells, imprinting spin textures varying along the azimuthal coordinate [67]. When a static magnetic field is applied in a plane, such as in the $x$-direction, individual spins start to precess around the $x$-axis. In Ref. [67], the periodic time evolution of $S_{z}$ reflecting the spatial structure of photon polarization has been measured by time- and spatially-resolved magneto-optical Kerr rotation. Therefore, combining spatially-resolved Kerr rotation with a static magnetic field may reveal the characteristic spin textures induced by vortex beams.
@@ Acknowledgements.
We thank Ken Morita for useful comments on the detection of in-plane spin textures. This work was supported by JST CREST (Grant No. JPMJCR24R5), JSPS KAKENHI (Grants No. JP23K20828, No. JP23K22492, No. JP25H00599, No. JP25H00609, No. JP25K07227, No. JP25K07198, and No. JP25K22011), a Grant-in-Aid for Transformative Research Areas (A) “Correlation Design Science” (Grant No. JP25H01250 and No. JP25H01251), “Evolution of Chiral Materials Science using Helical Light Fields” (Grant No. JP22H05131, JP23H04576 and JP25H01609), and “Chimera Quasiparticles for Novel Condensed-Matter Science” (Grant No. JP25H02112) from JSPS of Japan, and JSPS Program for Forming J-PEAKS (Grant No. JPJS00420230002).
@@ Appendix A Impurity vertex correction
Figure 7: Diagrammatic representation of ladder-type impurity scattering for vertex corrections ($\Lambda_{\mu}$) in spin-current response functions. The thick curves denote the Green’s function $G(k)$, and the dashed lines connected to the symbol $\times$” represent impurity scattering.
@@
In the main text, however, we consider the limit of weak impurity concentration, $\omega\tau\sim\Omega\tau\gg 1$. In this regime, the function, $\Pi_{00}$ can be evaluated by using $G^{\rm R}_{s}({\bm{k}},\varepsilon)=P\frac{1}{\varepsilon-E_{s}}-i\pi\delta(\varepsilon-E_{s})$ for $\Gamma\rightarrow 0$, where $P$ denotes the Cauchy principal part. Substituting this into Eq. (39), one finds ${\rm Re}\Pi_{00}=0$ and ${\rm Im}\Pi_{00}\propto\frac{1}{\omega}$. This indicates that in the limit of $\omega\tau\gg 1$, the spin vertex function reduces to the bare spin operator, $\hat{\Lambda}_{\mu}\approx\hat{S}_{\mu}$. While we have considered Rashba-type SOIs, the same conclusion is valid for ${\rm SU}(2)$ symmetric SOIs as well.